Solve the system using the substitution method. Show all your steps.
1. x+2y=6
3x-2y=2
3. 4x+y=7
2x+5y=-1
5. 3x+2y=-2
6x-y=6
Solve the system using the elimination method. Show all your steps.
7. -3x+3y=3
3x+y=9
9. -5x+12y=20
x-2y=-6
11. 3x+2y=1
4x+ 6y= 7
Solve the system using any algebraic method. Show all your steps.
13. 5x+ 7y= -2
2x-7y=9
14. x+3y=1
3x+7y=1
15. 4x+6y=8
2x+3y= 3
16. 8x-5y= -17
-2x+y=6
17. 3x-8y=0
-2x+5y=-2
18. 4x-6y=2
5x+3y= 1
19. 2x-5y=3
-4x+10y= -6
20. 8x+3y=10
-6x + y = -12
21. 5x+ 4y= -18
2x+3y=-24
22. CDs and Cassettes From 1990 to 1998, the manufacturer's shipments for audio cassettes A(in millions) and compact discs C(in millions) can be modeled by the equations
A= -31.8t + 322 Audio cassette shipments
C= 42.8t+110 Compact disc shipments
where t is the number of years since 1990. During what year did the number of compact discs shipped surpass the number of audio cassettes shipped?
Solve the system using any algebraic method. Show your steps.
30. 3x+7y= -1
2x +3y= 6
32. 3x-y= -2
5x+2y= 15
34. 2x + 3y= -6
3x- 4y= 25
36. 4x- 3y= 8
-8x +6y= 16
38. 7x+5y= -12
3x-4y=1

Solve the system. Then classify the system as consistent and independent, consistent and dependent, or inconsistent.
4, 4x+ 8y=8
x+ 2y=6
5. -5x+3y= -5
y= 5/3x + 1
6. x-2y=2
2x-y= -5
Solve the system using the substitution method. Show all steps.
7. 3x-y= -4
x+ 3y= -28
8. x+ 5y=1
-3x +4y= 16
9. 6x +y= -6
4x+3y= 17
Solve the system using the elmination method. Show all steps.
10. 2x- 3y= -1
2x+3y= -19
11. 3x-2y= 10
-6x +4y= -20
12. 2x +3y= 17
5x+8y= 20

Answer by checkley79(3341)

We don't normally do your homework.
We show you how to solve specific types of problems.

1. x+2y=6 or x=6-2y
3x-2y=2
3(6-2y)-2y=2
18-6y-2y=2
18-8y=2
-8y=-18+2
-8y=-16
y=-16/-8
y=2 ans.
x+2*2=6
x+4=6
x=6-4=2 ans.
Proof:
3*2-2*2=2
6-4=2
2=2
I leave the rest for you to practice with.

3. 4x+y=7
2x+5y=-1
5. 3x+2y=-2
6x-y=6
Solve the system using the elimination method. Show all your steps.
7. -3x+3y=3
3x+y=9 add.
---------------------------------
4y=12
y=12/4
y=3 ans.
3x+3=9
3x=9-3
3x=6
x=6/3
x=2 ans.
Proof:
-3*2+3*3=3
-6+9=3
3=3
I'll leave the rest for practice.
9. -5x+12y=20
x-2y=-6
11. 3x+2y=1
4x+ 6y= 7